# Images from disk and spherical accretions of Bardeen black hole surrounded by perfect fluid dark matter

### v1

@@ III Images of Bardeen black hole surrounded by PFDM illuminated by thin accretions disk
In this section, we use the ray-tracing method to first analyze the light ray propagation properties around the Bardeen black hole surrounded by PFDM, and then investigate the image of this black hole illuminated by a geometrically and optically thin accretion disk. We consider a static and spherically symmetric Bardeen black hole surrounded by PFDM, assuming that the accretion disk is located in the equatorial plane of the black hole and the observer is located at the North pole. Due to the strong gravitational lensing effect of the black hole, light rays emitted from the accretion disk may intersect the accretion disk multiple times before reaching the observer, significantly altering the optical appearance of the black hole. Therefore, we adopt the method proposed in Gralla:2019xty to first classify the light trajectories around the Bardeen black hole surrounded by PFDM, and then investigate the influence of dark matter parameter $\alpha$ and magnetic charge parameter $g$ on the light ray propagation around the black hole.
@@ III.1 Classification of the light rays
$\frac{dr}{d\phi}=\pm r^{2}\sqrt{\frac{1}{b^{2}}-V_{\mathrm{eff}}(r)},$ (9)
here, $\phi$ is the total change in azimuthal angle. Obviously, the equation of motion of the light ray is determined by the impact parameter $b$. One can define the number of light ray orbits $n=\phi/(2\pi)$. For $n=3/4$, the corresponding impact parameters are denoted as $b_{1}$ and $b_{4}$ (where $b_{1}<b_{4}$); for $n=5/4$, they are denoted as $b_{2}$ and $b_{3}$ (where $b_{2}<b_{3}$). Based on the number of times the light ray intersects the accretion disk, these light rays can be classified into three types Gralla:2019xty. (1)Direct emission: the light rays intersect the accretion disk only once, $1/4<n<3/4\Rightarrow b\in(0,b_{1})\cup(b_{4},+\infty)$. (2)Lensed ring emission: the light rays intersect the accretion disk twice, $3/4<n<5/4\Rightarrow b\in(b_{1},b_{2})\cup(b_{3},b_{4})$. (3) Photon ring emission: the light rays intersect the accretion disk at least three times, $n>5/4\Rightarrow b\in(b_{2},b_{3})$.
@@
In order to more intuitively illustrate the relationship between the number of photon orbits $n$ and the impact parameters, we plot the curves of the number of photon orbits as a function of the impact parameter in Fig.4(h) and present the light ray trajectories around the Bardeen black hole surrounded by PFDM. As the impact parameter increases, the photon orbit number $n$ gradually increases and reaches a peak at the critical impact parameter $b_{ph}$, and then gradually decreases. In addition, different values of dark matter parameters and magnetic charge parameters significantly affect the distribution width of the three types of emission light rays around the black hole. Therefore, we present the key characteristic parameters of the Bardeen black hole surrounded by PFDM in Table 1, including the event horizon, photon sphere radius, critical impact parameter with various dark matter parameters and magnetic charge parameters, as well as the corresponding impact parameter ranges for the three emission types of light rays. As discussed above, as the dark matter parameter $\left|\alpha\right|$ increases, the event horizon, photon sphere radius and critical impact parameter of the black hole all become larger, while the effect of the magnetic charge parameter $g$ is opposite to the dark matter parameter $\left|\alpha\right|$. Furthermore, as the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$ increase, the width of the lensed ring emission and the photon ring emission around the black hole will become wider. However, compared with the dark matter parameter $\alpha$, the effect of magnetic charge parameter $g$ on the width of lensed ring emission and photon ring emission is relatively smaller. Furthermore, the effect of magnetic charge on the optical appearance of black holes has also been extensively discussed He:2024amh; Guo:2023grt; He:2021htq; Akbarieh:2023kjv. Studies show that for the accretion images of static Bardeen black holes, the magnetic charge parameter $g$ decreases the size of the inner shadow and increases the total observed intensity in He:2021htq. Therefore, in the following studies, we will focus on investigating the influence of dark matter on the optical appearance of Bardeen black holes surrounded by PFDM.
@@
Table 1: The horizon radius $r_{h}$, photon sphere radius $r_{ph}$, critical impact parameter $b_{ph}$, the position of three emission types $b_{i}$ ($i=1,2,3,4$), photon ring width and lensed ring width with different coupling parameters. Here, we fix $M=1$.
@@ III.2 Observed intensities and optical appearances
Light rays emitted from the accretion disk may pass through the disk arbitrary times after being deflected by the black hole’s strong gravitational field. Each time it passes through the accretion disk, the light rays extract energy from it, significantly altering the optical appearance of the black hole. Due to the varying number of intersections between the three types of light rays and the accretion disk, they exhibit distinctly different observational signatures. Therefore, this section aims to investigate the effects of the three types of light on the optical appearance of the black hole and the observed images of the total intensity.
@@
here $r_{m}(b)$ is the transfer function, which describes the radial coordinate of the $m$-th intersection of the light ray with impact parameter $b$ with the accretion disk. The slope $dr/db$ is the scaling factor of the image. Therefore, a larger slope corresponds to strong demagnification and a smaller total luminosity. Subsequently, the transfer functions of the Bardeen black hole image for different dark matter parameter $\alpha$ cases are plotted in Fig.5. The black, gold, and red curves correspond to the transfer functions of direct emission, lensed ring emission, and photon ring emission, respectively. The slope of the transfer function of direct emission is almost $1$, which corresponds to the source profile after redshift and makes the main contribution to the total observed intensity. Compared with direct emission, the transfer function slope of the lensed ring emission is larger, which makes a bit contribution to the total photon flux, while the slope of the transfer function of the photon ring emission is much larger than $1$, and its contribution can be ignored. In addition, the increase of dark matter parameters $\left|\alpha\right|$ causes the first three transfer functions to move toward the direction of increasing impact parameter $b$, and the widths of the lensed ring and photon ring also significantly widen. This result is consistent with the data in Table 1.
@@
Figure 5: The first three transfer functions of a Bardeen black hole surrounded by PFDM with different magnetic charges $g$ and dark matter parameters $\alpha$. They represent the radial coordinates of the first (black), second (gold) and third (red) intersections with the emission disk. Here, we fix $M=1$.
@@
here, $I_{o}$ is the maximum intensity and the innermost stable circular orbit $r_{isco}$ can be obtained by Wang:2023vcv
@@
Next, we will investigate the images of Bardeen black hole surrounded by PFDM illuminated by two different accretion models. For a given emission function model, we can obtain the observed intensity of black hole via Eq.(12) and figure out the black hole image. For two different models, the results for the selected parameters are shown in Fig. 6 and Fig .7, where each figure can be divided into three columns. The black, gold, and red curves in the first column correspond to the observed intensities of direct, lensed ring, and photon ring emissions, respectively. The second column is the total observed intensity as a function of the impact parameters, and then project them into a two dimensional plane to explicitly present the optical appearance of the accretion disk in the third column. It is obvious from Fig. 6 and Fig .7 that the two different emission profile models of accretion disk lead to different black hole optical appearances. In the first emission model, the contributions of direct, lensed ring and photon ring emission are separated, which allows us to clearly observe a thin bright ring contributed by the photon ring emission. In contrast, in model II, the bright ring in the image is the result of the combined effects of these three types of emission, making it difficult to directly distinguish their individual contributions from the black hole’s optical appearance. This suggests that it is possible to infer the emission position and physical model of the accretion disk by analyzing the composition of the bright rings in astronomical observations. Of course, this requires more precise observation results in the future. In addition, a larger dark matter parameter $\left|\alpha\right|$ leads to a significant increase in the diameter of the inner shadow and bright ring of the black hole image. In addition, as the dark matter parameter $\left|\alpha\right|$ increases, the diameter of the inner shadow and bright ring of the black hole image significantly enlarges, while the total observed intensity decreases.
@@
Figure 6: Observational appearances of a thin disk with the first emission profile $I_{em1}(r)$ for different $\alpha$ with $g=0.1$, $M=1$. First column: the different observed intensities originated from the first (black), second (gold) and third (red) transfer function in Eq.12 respectively. Second column: the total observed intensities $I_{obs}/I_{o}$ as a function of impact parameter $b$. Third column: the optical appearance: the distribution of observed intensities into two-dimensional disks.
@@
Figure 7: Observational appearances of a thin disk with the third emission profile $I_{em2}(r)$ for different $\alpha$ with $g=0.1$, $M=1$.First column: the different observed intensities originated from the first (black), second (gold) and third (red) transfer function in Eq.12 respectively. Second column: the total observed intensities $I_{obs}/I_{o}$ as a function of impact parameter $b$.Third column: the optical appearance: the distribution of observed intensities into two-dimensional disks.
@@ IV Images of Bardeen black hole surrounded by PFDM illuminated by static spherical accretions
here, $\nu_{o}$ and $\nu_{e}$ denote the observed and emitted photon frequencies, respectively. The redshift factor is $g=\nu_{o}/\nu_{e}=f(r)^{1/2}$. $j_{e}(\nu_{e})$ is the emissivity per unit volume in the rest frame and usually taken the form $j_{e}(\nu_{e})\propto\delta(\nu_{r}-\nu_{e})/r^{2}$, here $\nu_{r}$ is the emitter’s rest-frame frequency. Additionally, $dl_{\text{prop}}$ is the infinitesimal proper length given by
@@
Furthermore, we investigated the observed intensity of Bardeen black holes surrounded by PFDM with different dark matter parameters $\alpha$ and depict them into a two dimensional plane, as shown in Fig. 8. In Fig. 8a, the peak in the total observed intensity at $b=b_{ph}$ manifests as the bright ring seen in the black hole image of Fig. 8b. For $b<b_{ph}$, the total observed intensities increases with the increase of the impact parameter $b$. On the contrary, when $b>b_{ph}$, a larger impact parameter $b$ corresponds to a smaller observed intensity. In addition, for larger dark matter parameters $\left|\alpha\right|$, the Bardeen black hole surrounded by PFDM exhibits a fainter bright ring and a larger central faint illuminating region in its image. Compared with the Bardeen black hole, the bright ring in the image of the Bardeen black hole surrounded by PFDM is fainter and the faint illuminating region in the center is larger, which is consistent with our previous conclusions.
@@
Figure 8: The total observed intensities (18) radiated from the static spherical accretion as a function of impact parameter, and the optical appearances of the Bardeen black hole surrounded by PFDM for selected parameters.
@@ V Conclusion and discussion
The imaging of black holes by the EHT collaboration provides a new approach to studying physical processes in strong fields regime and testing theories of gravity. Its observational data also offer an opportunity to test and constrain various theoretical models and gravitational theories. This paper investigates the shadow and accretion images of Bardeen black hole surrounded by the PFDM and constrains the dark matter parameters $\alpha$ using observational data from the EHT.
@@
As the preparation, we first investigate the influence of the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ on the fundamental characteristic quantities of the Bardeen black hole surrounded by the PFDM, including the event horizon radius $r_{h}$, photon sphere radius $r_{ph}$, critical impact parameter $b_{ph}$, and effective potential $V_{\text{eff}}$. The results show that as the dark matter parameter $\left|\alpha\right|$ increases, $r_{h}$, $r_{ph}$, and $b_{ph}$ all increase accordingly, while the increase of the magnetic charge parameter $g$ leads to these characteristic quantities to decrease. Furthermore, when $\left|\alpha\right|$ decreases, the peak of the effective potential rises and shifts in the direction of decreasing $r$. Conversely, an decrease in $g$ causes the peak of the effective potential to fall and shift in the direction of increasing $r$. Therefore, for the characteristic quantities of black holes $r_{h}$, $r_{ph}$, $b_{ph}$ and $V_{\text{eff}}$, there is a clear competing effect between the magnetic charge parameter $g$ and the dark matter parameter $\left|\alpha\right|$. In addition, we use the observational data of the supermassive black holes M87* and SgrA* to constrain the magnetic charge parameter $g$ and dark matter parameter $\alpha$ of the Bardeen black hole surrounded by PFDM. The magnetic charge parameter $g$ is consistent with observations throughout the entire parameter space ($g,\alpha$), while the dark matter parameter $\alpha$ is consistent with observations only when its value is above the red curve (see Fig. 3). Further analysis shows that the black hole SgrA* imposes stricter constraints on the dark matter parameter $\alpha$ than black hole M87*.
@@
Next, we analyze the images of the Bardeen black hole surrounded by PFDM illuminated by a geometrically and optically thin accretion disk. The optical appearance of the black hole shows significant differences for the two different emission profile models. In model I, the emissions of direct, lensed ring, and photon ring are separated from each other, and the thin ring produced by the photon ring emission can be clearly observed in the image. In model II, the emissions of the three are superimposed on each other to form the observed bright ring structure, and their respective contributions cannot be distinguished in the image. We look forward to discovering these differences through future higher-precision observational results, and also expect to use these differences to invert the emission profile model and location of the accretion disk. Furthermore, our results show that larger dark matter parameters $\left|\alpha\right|$ lead to an expansion of the inner shadow region of the black hole while reducing the observed intensity of the bright ring. Finally, we analyze the image of the Bardeen black hole surrounded by the PFDM by the illumination of a spherically symmetric accretion disk. Similar to the case of illumination from the thin accretion disk, larger dark matter parameters $\left|\alpha\right|$ increase the central faint illuminating region and reduce the brightness of the bright ring. Therefore, compared with the classical Bardeen black hole, the image of Bardeen black hole surrounded by the PFDM exhibits a larger inner shadow or faint illuminating region, while the bright ring appears fainter. It is worth noting that the images of other black holes surrounded by dark matter also exhibit a similar pattern: as the density of dark matter increases, the inner shadow in the black hole images enlarges Zeng:2025kqw; Luo:2025xjb; Li:2025ver.
@@
In conclusion, we constrain the dark matter parameters $\alpha$ using EHT observational data. Based on the investigation of different emission profile models, it is found that different models will significantly affect the image of the black hole. In the field of black hole observation, a series of achievements have been made, especially in accretion images and magnetic field deflection characteristics EventHorizonTelescope:2019dse; EventHorizonTelescope:2022wkp; EventHorizonTelescope:2025vum, but the understanding of the structure of black hole photon rings and the internal properties of accretion disks is still inadequate EventHorizonTelescope:2024whi; Johnson:2019ljv. Therefore, we expect that future higher-precision observation data will be able to observe these differences. Furthermore, our research demonstrates that dark matter parameters also significantly influence the black hole image, which provides a potential method to distinguish Bardeen black holes surrounded by PFDM from classical Bardeen black holes. We expect these preliminary results to provide some insights into using black hole shadows and images to study dark matter in the future.

### v2

@@ III Images of Bardeen black hole surrounded by PFDM illuminated by thin accretions disk
In this section, we use the ray-tracing method to first analyze the light ray propagation properties around the Bardeen black hole surrounded by PFDM, and then investigate the image of this black hole illuminated by a geometrically and optically thin accretion disk. We assume that the accretion disk is located in the equatorial plane of the black hole, and the observer is located at the North pole. Due to the strong gravitational lensing effect of the black hole, light rays emitted from the accretion disk may intersect the accretion disk multiple times before reaching the observer, significantly altering the optical appearance of the black hole. Therefore, we adopt the method proposed in Gralla:2019xty to first classify the light trajectories around the Bardeen black hole surrounded by PFDM, and then investigate the influence of dark matter parameter $\alpha$ and magnetic charge parameter $g$ on the light ray propagation around the black hole.
@@ III.1 Classification of the light rays
$\frac{dr}{d\phi}=\pm\frac{1}{r^{2}}\sqrt{\frac{1}{b^{2}}-V_{\mathrm{eff}}(r)}.$ (9)
Obviously, the trajectory of motion for the light ray is determined by the impact parameter $b$. One can define the number of light ray orbits $n=\phi/(2\pi)$, where $\phi$ is the total change in azimuthal angle. For $n=3/4$, the corresponding impact parameters are denoted as $b_{1}$ and $b_{4}$ (where $b_{1}<b_{4}$); for $n=5/4$, they are denoted as $b_{2}$ and $b_{3}$ (where $b_{2}<b_{3}$). Furthermore, based on the number of times the light ray intersects the accretion disk, light rays can be classified into three types Gralla:2019xty. (1)Direct emission: the light rays intersect the accretion disk at most once, $1/4<n<3/4\Rightarrow b\in(0,b_{1})\cup(b_{4},+\infty)$. (2)Lensed ring emission: the light rays intersect the accretion disk only twice, $3/4<n<5/4\Rightarrow b\in(b_{1},b_{2})\cup(b_{3},b_{4})$. (3) Photon ring emission: the light rays intersect the accretion disk at least three times, $n>5/4\Rightarrow b\in(b_{2},b_{3})$.
@@
In order to more intuitively illustrate the relationship between the number of photon orbits $n$ and the impact parameters, we plot the number of photon orbits as a function of the impact parameter in Fig.4 and present the light ray trajectories around the black hole. As the impact parameter increases, the photon orbit number $n$ gradually increases and reaches a peak at the critical impact parameter $b_{ph}$, and then gradually decreases. In addition, different values of dark matter parameters and magnetic charge parameters significantly affect the distribution width of the three types of emission light rays around the black hole. In Table 1, we present the key characteristic parameters of the Bardeen black hole surrounded by PFDM for various values of the dark matter and magnetic charge parameters. These include the event horizon, photon sphere radius, critical impact parameter, and the widths of the photon ring and lensed ring emissions. Obviously, as the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$ increase, the width of the lensed ring emission and the photon ring emission around the black hole will become wider. However, compared with the dark matter parameter $\alpha$, the effect of the magnetic charge parameter $g$ on the width of lensed ring emission and photon ring emission is relatively smaller. Moreover, the effect of magnetic charge on the optical appearance of black holes has also been extensively discussed He:2024amh; Guo:2023grt; He:2021htq; Akbarieh:2023kjv. The previous studies show that for the accretion images of static Bardeen black holes, the magnetic charge parameter $g$ decreases the size of the inner shadow and increases the total observed intensity in He:2021htq. Therefore, in the following studies, we will focus on the influence of dark matter parameter $\alpha$ on the optical appearance of Bardeen black holes surrounded by PFDM.
@@
Table 1: The event horizon $r_{h}$, photon sphere radius $r_{ph}$, critical impact parameter $b_{ph}$, the position of three emission types $b_{i}$ ($i=1,2,3,4$), photon ring width, and lensed ring width with different magnetic charges $g$ and dark matter parameters $\alpha$. Here, we fix $M=1$.
@@ III.2 Observed intensities and optical appearances
Light rays emitted from the accretion disk may pass through the disk arbitrary times after being deflected by the black hole’s strong gravitational field. Each time they pass through the accretion disk, the light rays extract energy from it, significantly contributing to the optical appearance of the black hole. Due to the varying number of intersections between the three types of light ray and the accretion disk, they exhibit distinctly different observational signatures. Therefore, this section aims to investigate the effects of the three types of light on the optical appearance of the black hole and the observed images of the total intensity.
@@
here, $r_{m}(b)$ is the transfer function, which describes the radial coordinate of the $m$-th intersection of the light ray with impact parameter $b$ with the accretion disk. The slope $dr/db$, being the scaling factor of the image, indicates that a larger value corresponds to strong demagnification and a smaller total luminosity. Subsequently, the transfer functions of the Bardeen black hole image for different dark matter parameters $\alpha$ are plotted in Fig.5. The black, gold, and red curves correspond to the transfer functions of direct emission, lensed ring emission, and photon ring emission, respectively. The slope of the transfer function of direct emission is almost $1$, which corresponds to the source profile after redshift and makes the main contribution to the total observed intensity. Compared with direct emission, the transfer function slope of the lensed ring emission is larger, which makes a bit of a contribution to the total photon flux. The slope of the transfer function of the photon ring emission is much larger than $1$, and its contribution can be ignored. In addition, the increase of dark matter parameters $\left|\alpha\right|$ causes the first three transfer functions to move toward the direction of increasing impact parameter $b$, and the widths of the lensed ring and photon ring also significantly widen. This result is consistent with the data in Table 1.
@@
Figure 5: The first three transfer functions of a Bardeen black hole surrounded by PFDM with different magnetic charges $g$ and dark matter parameters $\alpha$. They represent the radial coordinates of the first (black), second (gold), and third (red) intersections with the emission. Here, we fix $M=1$.
@@
here, $I_{o}$ is the maximum intensity, and the innermost stable circular orbit $r_{isco}$ can be obtained by Wang:2023vcv
@@
Next, we will investigate the images of Bardeen black hole surrounded by PFDM illuminated by two different accretion models. For a given emission function model, we can obtain the observed intensity of black hole via Eq.(12) and figure out the black hole image. The results for the selected parameters are shown in Fig.6 and Fig.7, where each figure can be divided into three columns. The black, gold, and red curves in the first column correspond to the observed intensities of direct, lensed ring, and photon ring emissions, respectively. The second column is the total observed intensity as a function of the impact parameter, and then project them into a two dimensional plane to explicitly present the optical appearance of the accretion disk in the third column. It is obvious from Fig.6 and Fig.7 that the two different emission models of accretion disk lead to different black hole optical appearances. In the first emission model, the contributions of direct, lensed ring, and photon ring emission are separated, which allows us to clearly observe a thin bright ring contributed by the photon ring emission. In contrast, in model II, the bright ring in the image is the result of the combined effects of these three types of emission, making it difficult to directly distinguish their individual contributions from the black hole’s optical appearance. This suggests that it is possible to infer the emission position and physical model of the accretion disk by analyzing the composition of the bright rings in astronomical observations. Of course, this requires more precise observation results in the future. In addition, as the dark matter parameter $\left|\alpha\right|$ increases, the diameter of the inner shadow and bright ring of the black hole image significantly enlarges, while the total observed intensity decreases.
@@
Figure 6: Observational appearances of a thin disk with the first emission profile $I_{em1}(r)$ for different $\alpha$ with $g=0.1$, $M=1$. First column: the different observed intensities originated from the first (black), second (gold), and third (red) transfer function in Eq.(12) respectively. Second column: the total observed intensities $I_{obs}/I_{o}$ as a function of impact parameter $b$. Third column: the optical appearance: the distribution of observed intensities into two-dimensional disks.
@@
Figure 7: Observational appearances of a thin disk with the third emission profile $I_{em2}(r)$ for different $\alpha$ with $g=0.1$, $M=1$.First column: the different observed intensities originated from the first (black), second (gold), and third (red) transfer function in Eq.(12) respectively. Second column: the total observed intensities $I_{obs}/I_{o}$ as a function of impact parameter $b$.Third column: the optical appearance: the distribution of observed intensities into two-dimensional disks.
@@ IV Images of Bardeen black hole surrounded by PFDM illuminated by static spherical accretions
here, $\nu_{o}$ and $\nu_{e}$ denote the observed and emitted photon frequencies, respectively. The redshift factor is $g=\nu_{o}/\nu_{e}=f(r)^{1/2}$. $j_{e}(\nu_{e})$ is the emissivity per unit volume in the rest frame and usually taken the form $j_{e}(\nu_{e})\propto\delta(\nu_{r}-\nu_{e})/r^{2}$, where $\nu_{r}$ is the emitter’s rest-frame frequency. Additionally, $dl_{\text{prop}}$ is the infinitesimal proper length given by
@@
Furthermore, we investigated the observed intensity of Bardeen black holes surrounded by PFDM with different dark matter parameters $\alpha$ and depicted them into a two dimensional plane, as shown in Fig.8. In Fig.8a, the peak in the total observed intensity at $b=b_{ph}$ manifests as the bright ring seen in the black hole image of Fig.8b. For $b<b_{ph}$, the total observed intensity increases with the increase of the impact parameter $b$. On the contrary, when $b>b_{ph}$, a larger impact parameter $b$ corresponds to a smaller observed intensity. In addition, for a larger dark matter parameter $\left|\alpha\right|$, the Bardeen black hole surrounded by PFDM exhibits a fainter bright ring and a larger central faint illuminating region in its image. Compared with the Bardeen black hole, the bright ring in the image of the Bardeen black hole surrounded by PFDM is fainter, and the faint illuminating region in the center is larger, which is consistent with our previous conclusions.
@@
Figure 8: The total observed intensity Eq.(18) radiated from the static spherical accretion as a function of impact parameter, and the optical appearances of the Bardeen black hole surrounded by PFDM for selected parameters.
@@ V Conclusion and discussion
The imaging of black holes by the EHT collaboration provides a new approach to studying physical processes in strong fields regime. Its observational data also offer an opportunity to constrain various theoretical models and theories of gravity. This paper investigates the shadow and accretion images of Bardeen black hole surrounded by PFDM and constrains the dark matter parameters $\alpha$.
@@
As preparation, we first investigate the influence of the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ on the fundamental characteristic quantities of the Bardeen black hole surrounded by PFDM, including the event horizon $r_{h}$, photon sphere radius $r_{ph}$, critical impact parameter $b_{ph}$, and effective potential $V_{\text{eff}}$. The results show that as the dark matter parameter $\left|\alpha\right|$ increases, $r_{h}$, $r_{ph}$, and $b_{ph}$ all increase accordingly, while the increase of the magnetic charge parameter $g$ leads to these characteristic quantities to decrease. Furthermore, when $\left|\alpha\right|$ decreases, the peak of the effective potential rises and shifts in the direction of decreasing $r$. Conversely, a decrease in $g$ causes the peak of the effective potential to fall and shift in the direction of increasing $r$. Therefore, for the characteristic quantities of black holes, $r_{h}$, $r_{ph}$, $b_{ph}$, and $V_{\text{eff}}$, there is a clear competing effect between the magnetic charge parameter $g$ and the dark matter parameter $\left|\alpha\right|$. In addition, we use the observational data of the supermassive black holes M87* and SgrA* to constrain the magnetic charge parameter $g$ and dark matter parameter $\alpha$ of the Bardeen black hole surrounded by PFDM. The magnetic charge parameter $g$ is consistent with observations throughout the entire parameter space ($g,\alpha$), while the dark matter parameter $\alpha$ is consistent with observations only when its value is above the red curve (see Fig. 3). Further analysis shows that the black hole SgrA* imposes stricter constraints on the dark matter parameter $\alpha$ than black hole M87*.
@@
Next, we analyze the images of the Bardeen black hole surrounded by PFDM illuminated by a geometrically and optically thin accretion disk. The optical appearance of the black hole shows significant differences for the two different emission profile models. In model I, the emissions of direct, lensed ring, and photon ring are separated from each other, and the thin ring produced by the photon ring emission can be clearly observed in the image. In model II, the emissions of the three are superimposed on each other to form the observed bright ring structure, and their respective contributions cannot be distinguished in the image. We look forward to discovering these differences through future higher-precision observational results, and also expect to use these differences to invert the emission profile model and location of the accretion disk. Furthermore, our results show that larger dark matter parameters $\left|\alpha\right|$ lead to an expansion of the inner shadow region of the black hole while reducing the observed intensity of the bright ring. Finally, we analyze the image of the Bardeen black hole surrounded by the PFDM by the illumination of a spherically symmetric accretion. Similar to the case of illumination from the thin accretion disk, larger dark matter parameters $\left|\alpha\right|$ increase the central faint illuminating region and reduce the brightness of the bright ring. Therefore, compared with the classical Bardeen black hole, the image of Bardeen black hole surrounded by PFDM exhibits a larger inner shadow or faint illuminating region. It is worth noting that the images of other black holes surrounded by dark matter also exhibit the same feature: as the density of dark matter increases, the inner shadow in the black hole images enlarges Zeng:2025kqw; Luo:2025xjb; Li:2025ver.
@@
In conclusion, we constrain the dark matter parameters $\alpha$ using EHT observational data. Based on the investigation of different emission profile models, it is found that different models will significantly affect the image of the black hole. Furthermore, our research demonstrates that dark matter parameters also significantly influence the black hole image, which provides a potential method to distinguish Bardeen black holes surrounded by PFDM from classical Bardeen black holes. Therefore, we expect that future higher-precision observation data will be able to observe these differences.
In fact, in the field of black hole observation, a series of achievements have been made, especially in accretion images and magnetic field deflection characteristics EventHorizonTelescope:2019dse; EventHorizonTelescope:2022wkp; EventHorizonTelescope:2025vum, but the understanding of the structure of black hole photon rings and the internal properties of accretion disks is still inadequate EventHorizonTelescope:2024whi; Johnson:2019ljv. We expect these preliminary results to provide some insights into using black hole shadows and images to study dark matter in the future.